OpenStudy (anonymous):

ABCD is congruent to EFGH, the measure of angle B = the measure of angle BCD = 90, and the measure of angle BAD = 131. What is the measure of angle GHE?

OpenStudy (anonymous):

in ABCD angleA = angleBAD = 131 deg angleB = 90 deg angleC = angleBCD = 90 deg since sum of all four interior angles of a quadrilateral is always 360 deg Hence angleD = 360 - (131 + 90 + 90) =360 - 311 = 49 deg since ABCD is congruent to EFGH, vertex D corresponds to vertex H and angleD = angleGHE so angleGHE = 49 deg

OpenStudy (anonymous):

thx!!

OpenStudy (anonymous):

Thanks for the help! :)

OpenStudy (anonymous):

Hey saffire86, can u pls become my fan as i need one more to complete 50 ?? o(( :})

OpenStudy (anonymous):

done :)

OpenStudy (anonymous):

Many thanks Saffire !!!!!! :D 50 more to go for a century !!!!